VEGAS Team Meeting

August 12, 2013


CALTECH, UC-B, & Columbia

Dan, Glenn, Mark,


Marty, Anish, John, Richard, Mark, Randy


  • Testing

Meeting Notes:

Mode #2 Testing:
  1. No data was taken in modes two or three during the weekend tests.
    • This was not a problem before the new HPC installations.
  2. One of the 1 hr long file with 10 msec integration there a utc jump in Bank B. This looks like a missing data. I think this test need to be repeated with the new HPC m/cs
  • Should retest with the new BOFs
  1. Freq axis in GBTIDL is not correct. This is a software issue.

ADC tests:
  1. New data has been received and is currently being plotted in an updated report. Glenn estimates that fine tuning of the ADCs could reduce spurs below the system threshold levels.

BOF files:
  1. Okay to continue DIBAS testing with existing BOF files, but they will be replaced when the ADC changes are incorporated.

Tuning executable threads to cores needs to occur
  1. need to find what cores OS uses and allocate around those cores.
  2. NRAO will look at how to accomplish this activity. (number of threads is < number of cores)

Other items:

Next Teleconference: August 17, 2013 .

-- MartyBloss - 2013-08-05
Topic revision: r1 - 2013-08-12, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback